Alice drives from Finley, Ohio to Richmond, Indiana to drop her sister off at school traveling at an average speed of 105 km/hr. On the way back she decides to drive a little slower at 95 km/hr. If her trip took a total of 8 hours, how far apart are the two towns?

Answer by Alan3354(69443)

Alice drives from Finley, Ohio to Richmond, Indiana to drop her sister off at school traveling at an average speed of 105 km/hr. On the way back she decides to drive a little slower at 95 km/hr. If her trip took a total of 8 hours, how far apart are the two towns?
----------------
Avg speed of the round trip = 2*105*95/(105+95) = 19950/200 = 99.75 km/hr
d = r*t = 99.75*8 = 798 km RT
--> 399 km each way
----------
PS I've been to Findlay, OH. Not to Finley, tho.